An air coref solenoid is of length 0.3 m , cross section area 1.2* 10-3 m2    nn hhahashas 2500 turns Around its central section a coil of 350 turns is wound The solenoid and the coil are electrically jnsulated from each other calculate the emf induced in the coil if the initial of 3A in the solenoid is reversed in 0.25 s 
 

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Magnetic field inside the solenoid intially=\mu_0nI.
Hence flux inside the solenoid=\mu_0nIA

After reversing the current, the direction of magnetic field also reverses. Hence change in magnetic field=\Delta B=2\mu_0nI.
The whole solenoid lies inside the coil and hence the magnetic flux inside the coil is ‘N’ timesthat inside solenoid.
Hence EMF induced in coil=\dfrac{\Delta \phi}{\Delta t}=\dfrac{2A\Delta B}{\Delta t}=\dfrac{2\mu_0NnI}{\Delta t}\dfrac{\Delta \phi}{\Delta t}=\dfrac{2AN\Delta B}{\Delta t}=\dfrac{2A\mu_0NnI}{\Delta t}